# DualMAR: Medical-Augmented Representation from Dual-Expertise Perspectives

### v1

@@ Appendix B KG Embeddings Comparison
Appendix C Details in Multilevel Attention
After generating $\mathbf{X}$ for all medical codes appears in the intersection of both training set $\mathcal{S}$ and concept vocabulary $\mathcal{C}$, our objective is to utilize EHR dataset to train both the model parameters and the embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{C}|\times d}$ for medical codes, and $d$ is the pre-defined embedding dimension for the previous GNN module. For the training process, we only focus on each patient $u$ with multiple admission records. An encoder is employed to transform all of their admission records into a patient representation vector $\mathbf{p}$, leveraging the embeddings in $\mathbf{X}$:
$\mathbf{p}=\text{Bi-Attention}(\mathbf{V}_{1},\mathbf{V}_{2},\ldots,\mathbf{V}_{T}|\mathbf{X}).$
In this study, “Bi-Attention” means that we implement a two-tier attention mechanism, detailed as follows:
a) Learning Admission Embeddings: Consider an admission record $\mathbf{V}_{\tau}$ containing $n$ medical codes. The embedding $\mathbf{x}_{i}$ for each medical code $c_{i}\in\mathbf{V}_{\tau}$ can be retrieved using $\mathbf{X}$, so that we can get code embeddings $\mathbf{E}^{\tau}_{V}$, which is also the subset of embeddings $\mathbf{E}$, for each admission. To aggregate these code embeddings into an admission embedding $\mathbf{v}_{\tau}$, we apply a global attention mechanism:
$\mathbf{E}^{\tau}_{V}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}]\in\mathbb{R}^{n\times m}$
$\mathbf{z}_{i}=\tanh(\mathbf{W}_{c}\mathbf{x}_{i})\in\mathbb{R}^{a}$
$\mathbf{v}_{\tau}=\sum_{i=1}^{n}\frac{\exp(\mathbf{z}_{i})}{\sum_{j=1}^{n}\exp(\mathbf{z}_{j})}\mathbf{x}_{i}\in\mathbb{R}^{m}.$
Here, $\mathbf{W}_{c}\in\mathbb{R}^{a\times m}$ is the weight matrix, where $a$ represent the attention dimension. The attention scores $\alpha{\tau}=[\alpha_{\tau}^{1},\alpha_{\tau}^{2},\ldots,\alpha_{\tau}^{n}]$ quantify the contribution of each medical code within an admission, and the weighted sum of $\mathbf{x}_{i}$ results in the admission embedding $\mathbf{v}_{\tau}$.
b) Learning Patient Embeddings: Once the embedding $\mathbf{v}_{\tau}$ for the $\tau$-th admission is computed, we derive the patient embedding by aggregating across all admissions. The admission embedding $\mathbf{v}_{\tau}$ is first projected into the patient embedding space through a simple linear classifier:
$\tilde{\mathbf{v}}_{\tau}=\sigma(\mathbf{W}_{u}\mathbf{v}_{\tau})\in\mathbb{R}^{p}.$
Here, $\mathbf{W}_{u}\in\mathbb{R}^{p\times m}$ is the corresponding weight matrix, and $p$ denotes the dimension of the patient embedding. The activation function is denoted as $\sigma$.
Assume patient $u$ has $T$ admission records as input features of model, we can get the admission embeddings $\mathbf{E}^{u}_{\mathbf{p}}$ after the code-level attention module. To aggregate $\mathbf{E}^{u}_{\mathbf{p}}$ into the final patient embedding $\mathbf{p}$, we then also apply the global attention across different admissions for each patient:
$\mathbf{E}^{u}_{\mathbf{p}}=[\tilde{\mathbf{v}}_{1},\tilde{\mathbf{v}}_{2},\dots,\tilde{\mathbf{v}}_{T}]\in\mathbb{R}^{T\times p}$
$\mathbf{r}_{\tau}=\tanh(\mathbf{W}_{v}\tilde{\mathbf{v}}_{\tau})\in\mathbb{R}^{b}$
$\mathbf{p}=\sum_{\tau=1}^{T}\frac{\exp(\mathbf{r}_{\tau})}{\sum_{\tau=1}^{T}\exp(\mathbf{r}_{\tau})}\tilde{\mathbf{v}}_{\tau}\in\mathbb{R}^{p}.$
In this case, $\mathbf{W}_{v}\in\mathbb{R}^{b\times p}$ is the weight matrix used for the transformation. The attention scores $\beta=[\beta_{1},\beta_{2},\ldots,\beta_{T}]$ capture the relative importance of each admission, and their weighted sum produces the final patient embedding $\mathbf{p}$.
Table 8: Statistics of MIMIC-III and MIMIC-IV datasets
Dataset MIMIC-III MIMIC-IV
# patients 7,493 10,000
Max. # visit 42 55
Avg. # visit 2.66 3.66
# codes 4,880 6,102
Max. # codes per visit 39 50
Avg. # codes per visit 13.06 13.38
Appendix D Experiment Implementation
D.1 Data Processing Details
To evaluate our proposed model, we focused on two well-established datasets in critical care research: MIMIC-III and MIMIC-IV. Both datasets are derived from de-identified clinical data collected at the Beth Israel Deaconess Medical Center in Boston, Massachusetts, and include detailed records from patients admitted to the Intensive Care Units (ICUs). MIMIC-III contains data from over 40,000 ICU admissions between 2001 and 2012, covering a wide range of information such as patient demographics, vital signs, laboratory results, diagnoses, and diagnostic codes. MIMIC-IV extends this dataset to approximately 60,000 ICU admissions from 2008 to 2019, reflecting more recent clinical practices and patient demographics.
To ensure distinct patient cohorts, we selected patients from MIMIC-IV who were admitted between 2018 and 2019, avoiding temporal overlap with MIMIC-III. We included only patients with multiple visits (# of visits $\geq$ 2) to ensure sufficient data for analysis. Specifically, for MIMIC-III, we divided the data into 6000 training, 500 validation, and 1000 testing samples. For MIMIC-IV, the distribution comprised 8000 training, 1000 validation, and 2000 testing samples. This approach allows us to leverage the longitudinal structure of EHR data, with the last visit of a patient serving as the label and all preceding visits as features. We aim to enhance the model’s ability to accurately predict critical care outcomes, contributing to advancements in predictive modeling and ultimately improving patient care through data-driven insights.
D.2 Parameter Settings
We consider node feature of graph learning module as the hierarchy KG embedding upon polar-space, which is $20002k=$ 2000 dimension vectors. Specifically, the embedding size $k=1000$ for both hierarchical and semantic representation in Diagnosis KG. We use two Graph Attention (GAT) Convolution layers for disease-lab graph, where the hidden dimension for both layers are 256, and we remain other settings as default values. The attention sizes $a=256$ and $b=256$ remain the same dimension as previous output, and to avoid over-fitting problem, we set dropout layers within both attention layers with $0.2$ dropout rate. Moreover, three three-level MLPs are adopted as decoders to help us get lab test embeddings with uniformly 256, 128 hidden dimensions, and we set dropout rate between layers are all 0.4 for either pretraining or direct training process. In addition, to finetune the lab test embeddings for various downstream tasks, a two-level classifiers with 256 hidden dimension and 0.5 dropout rate are used for transforming concatenated features to logits.
To retrieve precise hierarchical and semantic information from Diagnosis KG, we modify the source code of HAKE Zhang et al. [2020] and set training step as 180,000. For proxy-task learning, we use 10 and 10 epochs for joint and individual training respectively. We uniformly use Adam optimizer and decay learning rate for all training process, and we use 500 and 50 epochs for diagnosis prediction and heart failure prediction, respectively. Moreover, the initial learning rate in both pretraining and finetuning processes is 0.001. All programs are implemented using Python 3.10 and Pytorch 2.3.1 with CUDA 12.4 on a machine with two AMD EPYC 9254 24-Core Processors , 528GB RAM, and four Nvidia L40S GPUs.
Table 9: Prediction Results on MIMIC-IV for Diagnosis and HF Prediction.
Diagnosis Prediction HF Prediction
Models w-$\mathbf{F_{1}}$ R@10 R@20 AUC $\mathbf{F_{1}}$
Dipole 22.16(0.2) 36.21(0.2) 38.74(0.2) 84.80(0.3) 69.52(0.2)
Deepr 22.58(0.1) 36.79(0.1) 39.45(0.1) 83.61(0.5) 70.46(0.1)
RETAIN 23.11(0.8) 37.32(0.8) 40.15(0.6) 84.14(0.3) 71.23(0.2)
Timeline 23.76(0.2) 37.89(0.1) 40.87(0.1) 83.45(0.3) 72.30(0.2)
GRAM 24.39(0.1) 38.42(0.2) 41.62(0.3) 85.55(0.2) 69.82(0.4)
KAME 25.01(0.1) 38.86(0.2) 42.12(0.2) 84.80(0.2) 72.34(0.5)
CGL 25.74(0.2) 39.23(0.3) 42.67(0.2) 87.91(0.2) 70.71(0.3)
G-BERT 25.12(0.3) 39.91(0.2) 43.25(0.2) 85.76(0.2) 72.88(0.1)
HiTANet 24.53(0.2) 38.42(0.3) 41.89(0.1) 86.34(0.4) 71.35(0.2)
THAM 26.97(0.3) 43.07(0.4) 47.19(0.2) 87.43(0.3) 72.26(0.2)
DualMAR 29.87(0.3) 45.66(0.2) 51.73(0.3) 90.32(0.2) 73.54${}_{(\textbf{0.1})}$
Appendix E Details in Experiment
Table 8 shows the basic statistics for MIMIC-III and MIMIC-IV. Table 9 shows the result of main experiment which is the same setting as shown in MIMIC-III. Note that, we remain DualMAR in the same hyperparameter setting as model in MIMIC-III, but it still achieve superior performance across all baselines, which display the effectiveness and robustness of proposed model. Compared to MIMIC-III, MIMIC-IV has larger sample size and more longitude data, which accelerate the predictive performance of DualMAR within more complex scenarios, and we can tune the hyperparameters in model for pursuing even more accurate prediction for various downstream tasks.
@@ Appendix E Details in Experiment
Moreover, to evaluate the statistical significance of performance differences between our proposed DualMAR model and baseline models, we employed both the paired t-test and Wilcoxon signed-rank test. The t-test was chosen under the assumption of normally distributed performance metrics. To account for potential deviations from normality, the non-parametric Wilcoxon signed-rank test was also conducted. These tests allow us to robustly determine whether the observed improvements in DualMAR’s performance are statistically significant compared to the baselines.
@@
Here we only take the w-$F_{1}$ of Diagnosis Prediction in MIMIC-IV as an example to show the evaluation process on Table 10. We observe that the Wilcoxon signed-rank test consistently yielded p-values of $1.95\times 10^{-3}$, indicating significant differences between DualMAR and all baseline models. The paired t-tests further confirmed these results, with all p-values well below $1\times 10^{-10}$. The 95% confidence intervals for the t-tests showed a clear positive difference, reinforcing the superior performance of DualMAR in Diagnosis Prediction compared to each baseline. These results demonstrate the robustness of DualMAR’s improvements in weighted $F_{1}$ scores over the baselines.
@@
Table 10: Wilcoxon Signed-Rank Test and t-Test Results for DualMAR vs. Baseline Models with 10 runs. “p” means p-value for certain evaluation method, and “CI” means Confidence Interval calculated by T-testing method.
@@
DualMAR vs. Dipole 1.95e-03 1.62e-14 (7.41, 7.80)
DualMAR vs. Deepr 1.95e-03 2.01e-14 (7.17, 7.53)
DualMAR vs. RETAIN 1.95e-03 6.27e-10 (6.46, 7.39)
DualMAR vs. Timeline 1.95e-03 9.77e-13 (5.93, 6.39)
DualMAR vs. GRAM 1.95e-03 5.53e-13 (5.31, 5.67)
DualMAR vs. KAME 1.95e-03 1.23e-12 (4.87, 5.21)
DualMAR vs. CGL 1.95e-03 3.46e-11 (4.49, 4.88)
DualMAR vs. G-BERT 1.95e-03 7.65e-13 (5.62, 6.01)
DualMAR vs. HiTANet 1.95e-03 1.99e-12 (5.89, 6.27)
DualMAR vs. THAM 1.95e-03 4.57e-11 (3.98, 4.32)
@@
DualMAR-KGa 500 21.75(0.3) 33.27(0.4) 35.32(0.4)
DualMAR-KGa 2000 19.79(0.5) 32.12(0.4) 33.57(0.4)
DualMAR-KGb 500 23.87(0.3) 37.73(0.4) 39.18(0.4)
DualMAR-KGb 2000 24.58(0.3) 39.27(0.4) 40.32(0.4)
DualMAR-KGa+c 500 23.56(0.2) 37.43(0.3) 38.74(0.4)
DualMAR-KGa+c 2000 23.47(0.3) 37.89(0.4) 38.54(0.4)
DualMAR-KGb+c 500 24.32(0.3) 39.73(0.4) 40.15(0.4)
DualMAR-KGb+c 2000 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.2})}$ 41.86${}_{(\textbf{0.3})}$
Appendix F Details in Case Studies
@@
Table 12 shows the full results of both Rare-10 and Rare-20. It highlights DualMAR’s superior performance in predicting less frequent ICD codes across both subtasks, Rare-10 and Rare-20. Specifically, DualMAR consistently outperforms baseline models, achieving the highest Recall at all thresholds (R@5, R@8, R@15). Notably, for Rare-20, DualMAR significantly surpasses other models, particularly in R@8 and R@15, reflecting its strong capability in handling rare condition codes. These results underscore the effectiveness of DualMAR in scenarios where capturing rare but critical diagnoses is crucial.
Table 12 shows the full results of DualMAR with varying KGs and embedding sizes. Beyond analysis for w-$F_{1}$ in the case study, full table also demonstrates the Recall performance of DualMAR. The results indicate that incorporating KG components (KGb+c) consistently yields the highest Recall across R@10 and R@20, particularly with 2000-dimensional embeddings. Notably, KGb also shows strong performance, suggesting the significant contributions of fundamental Ontology-KG in enhancing diagnostic predictions. The comparison highlights the critical role of augmenting KG to achieve superior recall in diagnosis prediction tasks.

### v2

@@ 5.1 KG Embeddings Comparison
Most existing hierarchical KG embedding methods leverage Euclidean or non-Euclidean geometries to encode hierarchy. Euclidean-based approaches typically use coordinate projections to distinguish entity levels and retrieve hierarchical information more efficiently than hyperbolic or spherical methods, which often require additional steps or higher computational cost. To further improve efficiency, we decouple embedding training from graph learning by using link prediction tasks, allowing efficient GPU memory usage even on complex KGs. In our approach, each entity is projected onto polar coordinates to facilitate efficient embedding retrieval. As shown in Table 7, our polar-space embedding method, adapted from HAKE [44], outperforms five widely-used baselines [45, 47, 46, 48] in capturing both hierarchical and semantic features, evaluated by MRR and HITS@k metrics (k = 1, 3, 10).
5.2 The Significance of Improvement
@@
To evaluate the statistical significance of performance differences between our proposed DuaLK model and baseline models, we employed both the paired t-test and Wilcoxon signed-rank test. The t-test was chosen under the assumption of normally distributed performance metrics. To account for potential deviations from normality, the non-parametric Wilcoxon signed-rank test was also conducted. These tests allow us to robustly determine whether the observed improvements in DuaLK’s performance are statistically significant compared to the baselines.
@@
Here we only take the w-$F_{1}$ of Diagnosis Prediction in MIMIC-IV as an example to show the evaluation process on Table 8. We observe that the Wilcoxon signed-rank test consistently yielded p-values of $1.95\times 10^{-3}$, indicating significant differences between DuaLK and all baseline models. The paired t-tests further confirmed these results, with all p-values well below $1\times 10^{-10}$. The 95% confidence intervals for the t-tests showed a clear positive difference, reinforcing the superior performance of DuaLK in Diagnosis Prediction compared to each baseline. These results demonstrate the robustness of DuaLK’s improvements in weighted $F_{1}$ scores over the baselines.
@@
Table 8: Wilcoxon Signed-Rank Test and t-Test Results for DuaLK vs. Baseline Models with 10 runs. “p” means p-value for certain evaluation method, and “CI” means Confidence Interval calculated by T-testing method.
@@
DuaLK vs. RETAIN 1.95e-03 1.62e-14 (7.41, 7.80)
DuaLK vs. Timeline 1.95e-03 2.01e-14 (7.17, 7.53)
DuaLK vs. GCT 1.95e-03 6.27e-10 (6.46, 7.39)
DuaLK vs. CGL 1.95e-03 9.77e-13 (5.93, 6.39)
DuaLK vs. SeqCare 1.95e-03 5.53e-13 (5.31, 5.67)
DuaLK vs. RAM-EHR 1.95e-03 1.23e-12 (4.87, 5.21)
DuaLK vs. G-BERT 1.95e-03 3.46e-11 (4.49, 4.88)
DuaLK vs. HiTANet 1.95e-03 7.65e-13 (5.62, 6.01)
DuaLK vs. Med-BERT 1.95e-03 1.99e-12 (5.89, 6.27)
DuaLK vs. GT-BEHRT 1.95e-03 4.57e-11 (3.98, 4.32)
@@
DuaLK-KGa 500 21.75(0.3) 33.27(0.4) 35.32(0.4)
DuaLK-KGa 2000 19.79(0.5) 32.12(0.4) 33.57(0.4)
DuaLK-KGb 500 23.87(0.3) 37.73(0.4) 39.18(0.4)
DuaLK-KGb 2000 24.58(0.3) 39.27(0.4) 40.32(0.4)
DuaLK-KGa+c 500 23.56(0.2) 37.43(0.3) 38.74(0.4)
DuaLK-KGa+c 2000 23.47(0.3) 37.89(0.4) 38.54(0.4)
DuaLK-KGb+c 500 24.32(0.3) 39.73(0.4) 40.15(0.4)
DuaLK-KGb+c 2000 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.2})}$ 41.86${}_{(\textbf{0.3})}$
5.3 Details in Case Studies
@@
Table 9 shows the full results of both Rare-10 and Rare-20. It highlights DuaLK’s superior performance in predicting less frequent ICD codes across both subtasks, Rare-10 and Rare-20. Specifically, DuaLK consistently outperforms baseline models, achieving the highest Recall at all thresholds (R@5, R@8, R@15). Notably, for Rare-20, DuaLK significantly surpasses other models, particularly in R@8 and R@15, reflecting its strong capability in handling rare condition codes. These results underscore the effectiveness of DuaLK in scenarios where capturing rare but critical diagnoses is crucial.
Table 10 shows the full results of DuaLK with varying KGs and embedding sizes. Beyond analysis for w-$F_{1}$ in the case study, full table also demonstrates the Recall performance of DuaLK. The results indicate that incorporating KG components (KGb+c) consistently yields the highest Recall across R@10 and R@20, particularly with 2000-dimensional embeddings. Notably, KGb also shows strong performance, suggesting the significant contributions of fundamental Ontology-KG in enhancing diagnostic predictions. The comparison highlights the critical role of augmenting KG to achieve superior recall in diagnosis prediction tasks.
5.4 Prompting LLM
The following template in Table 11 is used to instruct LLMs to convert textual information into structured triples.
Name Prompt Template
Acknowledgment
This work is supported in part by the US National Science Foundation under grants 2047843 and 2437621. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.